OpenStudy (anonymous):

2x-y=2 2x-y=-2 I am thinking this is no solution or infinite solutions not sure

OpenStudy (anonymous):

do they have the same slope?

OpenStudy (anonymous):

2x-y=2 2x=2+y x=2+y/2 2(2+2y/2)-y=-2

OpenStudy (anonymous):

let me know if you need help finishing it

OpenStudy (anonymous):

do they have the same slope? (yes) and are they the same line? (no) therefor....

OpenStudy (anonymous):

Pretty much what I did was solved for x and substituted it into the second equation

OpenStudy (anonymous):

if they are the same line, there are an infinite number of solutions if they are parallel lines, there is no solution

OpenStudy (anonymous):

alright thats more the definition i needed thanks to Nismo45OZ and Philips13

OpenStudy (anonymous):

lines actually look parallel no there's no solution

OpenStudy (anonymous):

anytime